OpenStudy (anonymous):

PLEASE HELP ASAP mathematics will REWARD AND MEDAL 52.)2d-6=1/4d+1 53.)1/4x+5/8=x-1/8 54.)3/4z+5=1/4z-1 55.)2/3a-2/9=1/3a+7/9 56.)3/5-2/5b=4/5b

OpenStudy (nincompoop):

\[2d-6=\frac{ 1 }{ 4 }d+1\]

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

2d-1/4 d =7 so 7/4 d=7 d=4

OpenStudy (nincompoop):